OpenStudy (anonymous):

a car accelerates at a constant 2 miles/minute. if i accelerate for a distance of 10 miles, how fast am I going?

OpenStudy (amistre64):

i wonder if we integrate the accelration from 0 to 10 if we get the speed

OpenStudy (amistre64):

i know if we integrate velocity we get displacement ... im just unsure if hte same thing applies here :)

OpenStudy (amistre64):

\[\int 2 dt=2t\] so velocity is: v(t) = 2t but thats in time and not distance tho

OpenStudy (amistre64):

distance would be the integration of velocity \[\int 2t\ dt = t^2+t \]

OpenStudy (amistre64):

when t^2 + t = 10 we should have the time :)

OpenStudy (amistre64):

t^2+t=10 t^2+t+1/4 = 10 + 1/4 (t+1/2)^2 = 41/4 t = sqrt(41)/2 - 1/2 sounds resonable

OpenStudy (amistre64):

v((sqrt(41)-1)/2) = sqrt(41) - 1

OpenStudy (anonymous):

can't we use the formula s=ut+.5 a*t^2 so t=sqrt of 2s/a

OpenStudy (amistre64):

dunno, that looks more like a physics abomination :)

OpenStudy (anonymous):

yah, but it is applicable here as the acceleration is constant...

OpenStudy (amistre64):

what the "u" turn out to be? I see that .5*2 = 1 is the same as mine

OpenStudy (anonymous):

u means initial velocity here the body starts from rest so u=0..

OpenStudy (amistre64):

so im prolly off on my ints .... 2 2t t^2 + C ... yep